OpenStudy (anonymous):

log3(x) + log3(x-24) = 4 ??

OpenStudy (kamille):

well, you should know the rule \[\log _{a}b+\log _{a}c=\log _{a}(b*c)\] what can you do this this rule here?

OpenStudy (anonymous):

log3(x)*(x-24) = 4 --> log3(x^2-24x) = 4? :)

OpenStudy (kamille):

yeah, you are right! Also, you should know that \[\log _{a}b=c \] is the same as \[a ^{c}=b\] so now you get \[3^{4}=81\] and then \[\log _{3}81\] so you need to solve this equation \[\log _{3}(x ^{2}-24x)=\log _{3}81\] do you know how to do it?: ) p.s. do not forget to write and solve \[x ^{2}-24x>0\]

OpenStudy (anonymous):

I am still stuck lol! ): sorry... but can we get rid of the logs now? to make it x^24x = 81?

OpenStudy (kamille):

yes!

OpenStudy (anonymous):

Do I complete the square too? thank you so much! :o)

OpenStudy (kamille):

yeah, and then solve \[x ^{2}-24x>0\] to check if your solution is right